# A Sloppy approach to QSP: XAI enabling fit-for-purpose models

### v2

@@ 2 Results
2.1 A Sloppy QSP Workflow
A primary outcome of this work is the formalization of a QSP workflow based on principles and techniques of “Sloppy modeling.” As summarized briefly in the introduction, our proposed workflow is illustrated graphically in Fig 1. Here we elaborate on this process before presenting the results of the validation studies.
In our approach, the quality of a mathematical model is always measured against its ability to facilitate answering specific scientific questions, illustrated by the box “QSP Objectives” in the upper left. In QSP, the ultimate objectives are outcomes such as target discovery, toxicology, personalized medicine, or other similar objectives. The model realizes these high-level objectives through concrete calculations (box in upper right), such as testing a hypothesis by direct comparison against data, uncertainty quantification or virtual populations, designing experiments for further study, etc. To develop a model fit for this purpose, we propose the workflow summarized in the flow diagram between these two boxes.
In the first step, a detailed model is constructed by gathering mechanistic structures from the available literature, resulting in what we term a literature-derived model. In traditional practice, model building is as much an art as a science, relying heavily on the modeler’s physical and mathematical intuition to select structures that seem most relevant. Yet in complex biological systems, intuition is slow to develop: assembling an initial model may take months, while true insight into system behavior often requires years or even decades. Crucial modeling decisions—such as which mechanisms to include or omit—are shaped by expert judgment, but this judgment is inherently vulnerable to gaps in knowledge, disciplinary biases, and cognitive heuristics. By contrast, the literature-derived model offers epistemic neutrality: it defers structural choices to the collective body of published research, ensuring that no single theoretical perspective dominates the model’s architecture. In highly complex systems where causal pathways remain incomplete or ambiguous, this strategy guards against overemphasizing the familiar while overlooking the novel or counterintuitive [68, 69]. Moreover, by deliberately including all documented mechanisms upfront, our approach streamlines the modeling process: it reduces the time and expertise required for model construction, while anticipating the later application of automated, context-specific model reduction. Rather than relying on human intuition to guess which mechanisms matter, the reduction step systematically identifies and preserves the structures essential to the phenomena of interest. Literature-derived models can also be reused for later applications. This shift standardizes the model building workflow, making it more automated, transparent, and objective, rendering it especially well-suited for scalable QSP applications.
@@ 2.1 A Sloppy QSP Workflow
At this point, it is possible that the model cannot fit the data very well. If that is the case, there are likely discrepancies between the mechanisms known in the literature and those of the real system. If the fit is inadequate, the model could be further refined by adding additional terms to the model. This was not necessary for any of the models used in this study; however, there are indications in our fits that models could be further improved in future work. When a model fails to capture a target phenomenon, it does so transparently: not because it was skewed by human judgment, but because the requisite knowledge has not yet been formally articulated. Thus, even the failure of a literature-derived model is highly informative.
Generally, we do not expect the parameters in the calibrated model to be tightly constrained by the available system data. The theory of sloppy models demonstrates that QSP model parameters are generically unidentifiable. However, a few combinations of parameters are well-constrained after fitting the data. These are known as the stiff parameters in the sloppy model literature, and these are the parameters that encode the information the model has taken from the system data.
@@
To circumvent these challenges, we apply a parameter reduction algorithm known as the Manifold Boundary Approximation Method (MBAM)[32]. This algorithm builds reduced-order models that include only parameters relevant to the given queries. This reduced model is formally a special case of the full model; however, it is projected onto the relevant parameters that actively participate in the propagation of information from data to queries of interest. To borrow the terminology of [70], we apply Occam’s razor, but take care that it does not cut so deep the model loses its ability to “reach” behaviors for which we don’t have experimental data. The query paradigm preserves parameters that are important not only for fitting available data, but also for generating reliable predictions, thus preserving relevant mechanisms even if they were unobserved experimentally. We propose conducting the target QSP calculations using this reduced model to avoid the problem discussed above. In the following sections, we demonstrate how these reduced models reveal the relevant control mechanisms underlying a behavior, are predictive under novel experimental conditions, and enable stable propagation of uncertainty while facilitating reasoning about mechanisms underlying different responses.
@@ 2.2 Coagulation Modeling
The coagulation cascade is a particularly well-studied system with characteristics that present typical challenges to QSP models, all of which are addressed by our workflow. In this cascade, Factors VII (stable factor) and XII (Hageman factor) are activated by tissue damage, triggering two separate cascades which both activate Factor II (prothrombin), which in turn activates the clotting factor fibrin. The underlying biology is well-characterized in the literature, with several previous studies having produced extensive literature-derived models. The system spans multiple biological scales, from molecular (e.g., enzyme kinetics) to larger-scale physiological outcomes (e.g., clot formation). It also exhibits strong nonlinearities and feedback mechanisms, which pose challenges for parameter identification, calibration, and sensitivity analysis. On the experimental side, there is a wealth of in vitro, in vivo, and clinical data available for rigorous model validation and comparison studies. Furthermore, many therapeutic agents, particularly anticoagulants, target the coagulation pathway, making the system relevant to realistic drug development scenarios. In particular, models of the coagulation cascade are known to suffer from parameter unidentifiability and have been used as validation studies for QSP workflows [25, 71, 12, 72].
@@ 2.2.1 Minimal Model
The parameter values reported by Nayak et al. give us a calibrated model; however, sensitivity analysis reveals that the model is sloppy, and most of these parameters are unidentifiable from these data. We wish to explore the ability of a reduced-order model to make predictions for QSP-specific queries. The queries we chose are thrombin time series (Factor IIa) in response to various doses of Factor VIIa and Factor Xa added to both normal and Factor VII-deficient human plasma, for a total of 28 time series. These are the same time series used in model calibration. The thrombin response in each of these time series is adaptive[78, 33], i.e., after an initial response to the dosage, it returns to baseline levels after long times. Our final coarse model based on these queries has five parameters, as is represented graphically in the foreground of Fig 2, along with a typical Thrombin profile for both the full and reduced models.
Notice how the reduced model dramatically highlights the key mechanism driving the adaptive behavior of the system. In response to an initial dose, Factor X is activated, which then catalyzes the activation of Factor II. This initial catalysis is relatively weak, but it subsequently triggers a feed-forward loop in which IIa activates Factor V, which then forms a complex with Xa. The Va-Xa complex is a much stronger catalyst for activation of Factor II and is responsible for the bulk of the activation in each time series. This causal narrative is consistent with how QSP scientists qualitatively understand the coagulation process. Our pipeline algorithmically finds a minimal mathematical explanation from the literature-derived model that encapsulates the human intuition in a quantitative model. The next section demonstrates the ability of the reduced model to answer key QSP questions.
@@ 2.2.2 Predictive Power of the Minimal Model
Fig 3: Predictions of the reduced order coagulation model compared to data. The mechanism of action (incoherent feed-forward loop) guarantees that the max thrombin level is inversely proportional to the max time to achieve maximum thrombin activation (Left). Consequently, the area under the curve (right) is approximately constant over a wide range of control parameters. The model accurately predicts this tradeoff, even when only one Thrombin profile is used as a data query.
@@
The reduced-order models we present here offer several advantages over the full, literature-derived model for facilitating QSP studies. First, they provide human-interpretable insights into the key mechanisms responsible for driving the phenomena of interest. Beyond the biochemical simplifications evident from the reduced network structure, the fitted parameter values further highlight the relative importance of specific reactions in shaping the model’s predictions. Second, the reduced models are computationally more efficient and numerically more stable than the full models. In our study, direct simulation of the SBML-extracted full model revealed extreme sensitivity to numerical details: different integration algorithms produced wildly divergent solution profiles. Only by evaluating multiple solvers at extremely high precision could we determine which solution accurately reflected the underlying mathematical equations. By contrast, the reduced-order model achieved a two orders of magnitude reduction in computation time while simultaneously eliminating sensitivity to solver choice, enhancing numerical stability, and increasing overall confidence in model-based conclusions. Notably, the process of addressing parameter identifiability through model reduction also resolves many of the broader challenges associated with model complexity. We now illustrate these advantages in a case study on HIV infection dynamics, showcasing the full sloppy modeling workflow from start to finish.
@@ 2.3.1 Literature-Derived and Calibrated Models
Note that previous studies used the terms “controller” and “non-controller” to indicate whether a specific monkey responded to treatment. We instead use the terms “responder” and “nonresponder” to avoid confusion with the untreated comparison group, i.e., the statistical control. Referring to all groups other than the control group using variants of the word “control” undermines the clarity of terminology. Furthermore, “elite controllers” traditionally refers to individuals able to resist HIV without intervention[86, 87, 88]. In this work, we do not distinguish between innate resistance and resistance acquired through the bNAb intervention.
@@
Fig 5: Model fits to the aggregate untreated data (top left), responder DEWL data (top right), and two fits to the nonresponder DF06 data showing a good fit (bottom left) and an inferior but statistically significant fit that activates different mechanisms (bottom right).
@@ 2.3.2 Reduced Models and Mechanistic Insights
Although the pharmacokinetic predictions for both drugs are retained, the Hill function enhancing viral clearance is simplified to depend largely on $A_{1}$ for therapeutic benefit. Since the drugs are modeled to have symmetric effects, the presence of only a single drug is mathematically sufficient to predict the treated response. A more detailed model that distinguishes between the two drugs’ mechanisms of action may provide more information about their synergy or competition.
@@ 2.3.3 Drug target identification
2.3.4 Uncertainty Quantification, Virtual Populations, and Optimal Experimental Design
One of the advantages of mathematical modeling is the potential for rigorous Uncertainty Quantification (UQ) for downstream applications of interest. In QSP, this is especially important for assessing risk, designing new experiments, or fulfilling other objectives. When a model is unidentifiable, however, UQ methods are ill-conditioned and numerically unstable. In this section we demonstrate another application of our workflow by performing rigorous Bayesian UQ and virtual population generation on a reduced order model. These techniques inform both development and regulatory agencies about potential responses and risks in the broader population.
We ask: given a typical response to an untreated viral infection, can one quantify the range of the response to treatment in a broader population? We take as our starting calibrated model the parameter values fit to the aggregated, untreated data (Fig 4a). The queries are the viral load predictions, both with and without bNAb therapy. The time series without therapy is the data query and the prediction under treatment conditions are the target predictions. Our ultimate goal is to quantify uncertainty in these downstream predictions and generate a virtual population of potential responses. Finally, we consider how experimental design can reduce uncertainty when predicting an individual’s response in a personalized medicine scenario.
@@ 2.3.4 Uncertainty Quantification, Virtual Populations, and Optimal Experimental Design
MBAM gives a 10-parameter reduced model for these queries (Eq. (12) in the methods section). Since these queries are a subset of those used in section 2.3.2, the model is even simpler than before. Inspecting the resulting equations (see methods), six of the remaining parameters characterize the bNAb dynamics and therefore are trivially structurally unidentifiable from the untreated profile. For now, we hold these at their nominal values and consider the uncertainty in the remaining four parameters. We then used this reduced model to perform uncertainty quantification using two UQ techniques: the delta method[93] and Bayesian posterior sampling (see methods for more details). Fig 7 summarizes the results of the UQ analysis for both methods for predictions under untreated and treated scenarios.
Fig 7: Uncertainty quantification of a Virtual Population for untreated (first column) and treated (second column) predictions using the delta method (first row) and Bayesian sampling (second row). Uncertainty in the untreated predictions is reasonably small where the model interpolates data. Uncertainty in the treated predictions, however, is quite large due to parameters that are not well-constrained by the data. The Bayesian ensemble of predictions gives more reasonable UQ estimates than the delta method, which relies on model linearization. For both methods, uncertainty is still too large to predict clinical outcome.
First, consider the uncertainty for untreated profiles, shown in the first column of Fig 7, for the delta and Bayesian methods, respectively. Because the model was calibrated to untreated viral load measurements, the uncertainty is relatively small for predictions near the fitted data. Uncertainty is larger at early times, before any measurements are available. Notice the good agreement between the uncertainty in these predictions for the two methods.
The uncertainty for the treated profiles are shown in the second column of Fig 7. The uncertainty in these predictions is very large for both methods, indicating the information in the untreated response is not sufficient to constrain predictions under treatment scenarios. This uncertainty represents the potential range of responses represents that could be exhibited by individuals in a population.
This result helps to dispel a potential confusion about the role of parameter reduction in our methodology. The objective of the parameter reduction step is not to artificially reduce uncertainty in predictions by removing unidentifiable parameters. Rather, it facilitates quantifying the informational relationship between training data and predictions. We see here that large uncertainties can arise if parameters are relevant for predictions while remaining unconstrained by data. This uncertainty is particularly relevant to pharmaceutical development as these potential responses inform risk assessment and regulatory review.
The uncertainty in the treated profiles also provides a useful comparison of the delta and Bayesian methods. While both methods agree that uncertainty in these predictions is large, the delta method predicts unrealistically large uncertainties that are symmetrically distributed around the nominal predictions. These features arise because the delta method is based on a linearization of the model and fails to capture saturation effects associated with large parameter uncertainties. In contrast, the Bayesian sampling fully incorporates the nonlinear effects of the model. In a typical QSP model, rigorously sampling the full Bayesian posterior is almost always infeasible due to the need to sample the large, sloppy subspace. In contrast, the posterior presented here was sampled in about 3 hours using a laptop with an Intel /Core i7 3Ghz processor with 4 cores and 16GB of RAM.
Although the Bayesian posterior sampling is more rigorous, both methods of uncertainty quantification demonstrated that the treated predictions are not accurate and should not be used to make a clinical judgment for specific patients. The reduced model, using best-fit parameters, does predict a successful treatment response, with the viral load dropping exponentially. In the absence of uncertainty quantification to complement these predictions, this result could be erroneously interpreted as signaling successful treatment, when in fact no patient-specific conclusions should be drawn without further information. Such UQ calculations are critical for assessing risk in regulatory decision-making.
The simplicity of our reduced model also allows us to understand the source of uncertainty in the model predictions, which can inform experimental design and personalized medicine. Figure 8 shows the pair plot histograms of the Bayesian posterior sample. Parameters $\theta_{1}$ and $\theta_{2}$, which determine the initial concentrations of T-cells and free virus, have large standard deviations and are the primary source of uncertainty in untreated predictions at early times. Furthermore, $\theta_{1}$ and $\theta_{2}$ are also highly correlated, suggesting that their ratio is well-constrained. Since the untreated predictions become constrained at later times, only the ratio $\frac{\theta_{1}}{\theta_{2}}$ is needed to make accurate predictions for untreated scenarios. For treated predictions, the influence of drug concentration $A$ is modeled by an enhanced viral clearance rate of $\theta_{2}+A$. Since $\theta_{2}$ is unconstrained and could be large or small, it is impossible to predict whether the drug leads to a perceptible increase in clearance without additional data. If additional measurements of viral load at early times were acquired and used to recalibrate the model, $\theta_{2}$ could be better constrained, and uncertainty in the treated predictions could be reduced. Such additional information would constrain the clinical outcome in a personalized medicine scenario, which is not possible with the current data set.
Fig 8: Pair plot for the Bayesian posterior Virtual Population. The primary source of uncertainty comes from parameters 1 and 2, which are largely unconstrained, but highly correlated.

### v3

@@ 2 Results
2.1 A CRISP QSP Workflow
A primary outcome of this work is the formalization of a QSP workflow based on principles and techniques of Information Geometry. As summarized briefly in the introduction, the Contextualized Reduction for Identifiability and Scientific Precision (CRISP) workflow is illustrated graphically in Fig 1. Here we elaborate on this process before presenting the results of the validation studies.
In our approach, the quality of a mathematical model is always measured against its ability to facilitate answering specific scientific questions, illustrated by the box “QSP Objectives” in the upper left. In QSP, the ultimate objectives are outcomes such as target discovery, toxicology, personalized medicine, and so forth. The model realizes these high-level objectives through concrete calculations (box in upper right), such as testing a hypothesis by direct comparison against data, uncertainty quantification or virtual populations, designing experiments for further study, etc. To develop a model fit for this purpose, we propose the workflow summarized in the flow diagram between these two boxes.
In the first step, a detailed model is constructed by gathering mechanistic structures from the available literature, resulting in what we term a literature-derived model. In traditional practice, model building is as much an art as a science, relying heavily on the modeler’s physical and mathematical intuition to select structures that seem most relevant. Yet in complex biological systems, intuition is slow to develop: assembling an initial model may take months, while true insight into system behavior often requires years or even decades. Crucial modeling decisions—such as which mechanisms to include or omit—are shaped by expert judgment, but this judgment is inherently vulnerable to gaps in knowledge, disciplinary biases, and cognitive heuristics. By contrast, the literature-derived model offers epistemic neutrality: it defers structural choices to the collective body of published research, ensuring that no single theoretical perspective dominates the model’s architecture. In highly complex systems where causal pathways remain incomplete or ambiguous, this strategy guards against the universal human tendency to overemphasize the familiar while overlooking the novel or counterintuitive [68, 69]. Moreover, by deliberately including all documented mechanisms upfront, our approach streamlines the modeling process: it reduces the time and expertise required for model construction, while anticipating the later application of automated, context-specific model reduction. Rather than relying on human intuition to guess which mechanisms matter, the reduction step systematically identifies and preserves the structures essential to the phenomena of interest. Literature-derived models can also be reused for later applications. This shift standardizes the model building workflow, making it more automated, transparent, and objective, rendering it especially well-suited for scalable QSP applications.
@@ 2.1 A CRISP QSP Workflow
At this point, it is possible that the model cannot fit the data very well. If that is the case, there are likely discrepancies between the mechanisms known in the literature and those of the real system. If the fit is inadequate, the model could be further refined by adding additional terms to the model. This was not necessary for any of the models used in this study; however, there are indications in our fits that models could be further improved in future work. When a model fails to capture a target phenomenon, it does so transparently: not because it was skewed by human judgment, but because the requisite knowledge has not yet been formally articulated. Thus, even the failure of a literature-derived model is highly informative, as it identifies an area where more fundamental science can be done.
Generally, we do not expect the parameters in the calibrated model to be tightly constrained by the available system data. Results from Information Geometry demonstrate that QSP model parameters are generically unidentifiable. However, a few combinations of parameters are well-constrained after fitting the data. These are known as the stiff or relevant parameters, and these are the parameters that encode the information the model has taken from the system data.
@@
To circumvent these challenges, we apply a parameter reduction algorithm known as the Manifold Boundary Approximation Method (MBAM)[32]. This algorithm carves reduced-order models from the literature-derived model, which include only parameters relevant to the given queries. This reduced model is formally a special case of the full model; however, it is projected onto the relevant parameters that actively participate in the propagation of information from data to queries of interest. To borrow the terminology of [70], we apply Occam’s razor, but take care that it does not cut so deep the model loses its ability to “reach” behaviors for which we don’t have experimental data [71]. The query paradigm preserves parameters that are important not only for fitting available data, but also for generating reliable predictions, thus preserving relevant mechanisms even if they were unobserved experimentally. We propose conducting the target QSP calculations using this reduced model to avoid the problem discussed above. In the following sections, we demonstrate how these reduced models reveal the relevant control mechanisms underlying a behavior, are predictive under novel experimental conditions, and enable stable propagation of uncertainty while facilitating reasoning about mechanisms underlying different responses.
@@ 2.2 Coagulation Modeling
The coagulation cascade is a particularly well-studied system with characteristics that present typical challenges to QSP models, all of which are addressed by our workflow. In this cascade, Factors VII (stable factor) and XII (Hageman factor) are activated by tissue damage, triggering two separate cascades which both activate Factor II (prothrombin), which in turn activates the clotting factor fibrin. The underlying biology is well-characterized in the literature, with several previous studies having produced extensive literature-derived models. It also exhibits many of the challenges and opportunities typical of QSP models. The system spans multiple biological scales, from molecular (e.g., enzyme kinetics) to larger-scale physiological outcomes (e.g., clot formation). It also exhibits strong nonlinearities and feedback mechanisms, which pose challenges for parameter identification, calibration, and sensitivity analysis. On the experimental side, there is a wealth of in vitro, in vivo, and clinical data available for rigorous model validation and comparison studies. Furthermore, many therapeutic agents, particularly anticoagulants, target the coagulation pathway, making the system relevant to realistic drug development scenarios. In particular, models of the coagulation cascade are known to suffer from over-parameterization and have been used as validation studies for QSP workflows [25, 72, 12, 73].
@@ 2.2.1 Minimal Model
The parameter values reported by Nayak et al. give us a calibrated model; however, sensitivity analysis reveals that most of the parameters are unidentifiable from these data. We wish to explore the ability of a reduced-order model to make predictions for QSP-specific queries. The queries we chose are thrombin time series (Factor IIa) in response to various doses of Factor VIIa and Factor Xa added to both normal and Factor VII-deficient human plasma, for a total of 28 time series. These are the same time series used in model calibration. The thrombin response in each of these time series is adaptive[79, 33], i.e., after an initial response to the dosage, it returns to baseline levels after long times. Our final coarse model based on these queries has five parameters, as is represented graphically in the foreground of Fig 2, along with a typical thrombin profile for both the full and reduced models.
Notice how the reduced model dramatically highlights the key mechanism driving the adaptive behavior of the system. In response to an initial dose, Factor X is activated, which then catalyzes the activation of Factor II. This initial catalysis is relatively weak, but it subsequently triggers a feed-forward loop in which IIa activates Factor V, which then forms a complex with Xa. The Va-Xa complex is a much stronger catalyst for activation of Factor II and is responsible for the bulk of the activation in each time series. This causal narrative is consistent with how QSP scientists qualitatively understand the coagulation process. Our pipeline algorithmically finds a minimal mathematical explanation from the literature-derived model that encapsulates the human intuition in a quantitative model with only five dynamical variables and four parameters. The next section demonstrates the ability of the reduced model to answer key QSP questions.
@@ 2.2.2 Predictive Power of the Minimal Model
Fig 3: Predictions of the reduced order coagulation model compared to data. The mechanism of action (incoherent feed-forward loop) guarantees that the max thrombin level is inversely proportional to the max time to achieve maximum thrombin activation (a). Consequently, the area under the curve (b) is approximately constant over a wide range of control parameters. The model accurately predicts this trade-off, even when only one thrombin profile is used as a data query (a 94% reduction in data demand).
@@
The reduced-order models we present here offer several advantages over the full, literature-derived model for facilitating QSP studies. First, they provide human-interpretable insights into the key mechanisms responsible for driving the phenomena of interest. Beyond the biochemical simplifications evident from the reduced network structure, the fitted parameter values further highlight the relative importance of specific reactions in shaping the model’s predictions. Second, the reduced models are computationally more efficient and numerically more stable than the full models. In our study, direct simulation of the SBML-extracted full model revealed extreme sensitivity to numerical details: different integration algorithms produced wildly divergent solution profiles. Only by evaluating multiple solvers at extremely high precision could we determine which solution accurately reflected the underlying mathematical equations. By contrast, the reduced-order model achieved a two orders of magnitude reduction in computation time while simultaneously eliminating sensitivity to solver choice, enhancing numerical stability, and increasing overall confidence in model-based conclusions. Notably, the process of addressing parameter identifiability through model reduction also resolves many of the broader challenges associated with model complexity. We now illustrate these advantages in a case study on HIV infection dynamics, showcasing the full CRISP workflow from start to finish.
@@ 2.3.1 Literature-Derived and Calibrated Models
Note that previous studies used the terms “controller” and “non-controller” to indicate whether a specific monkey responded to treatment. We instead use the terms “responder” and “nonresponder” to avoid confusion with the untreated comparison group, i.e., the statistical control. Referring to all groups other than the control group using variants of the word “control” undermines the clarity of terminology. Furthermore, “elite controllers” traditionally refers to individuals able to resist HIV without intervention[87, 88, 89, 90]. For the purposes of this study, we do not wish to distinguish between innate resistance and resistance acquired through the bNAb intervention.
@@
Fig 5: Model fits to the aggregate untreated data (A), responder DEWL data (B), and two fits to the nonresponder DF06 data showing a good fit (C) and an inferior but statistically significant fit that activates different mechanisms (D).
@@ 2.3.2 Reduced Models and Mechanistic Insights
Although the pharmacokinetic predictions for both drugs are retained, the Hill function enhancing viral clearance is simplified to depend largely on $A_{2}$ for therapeutic benefit. Since the drugs are modeled to have symmetric effects, the presence of only a single drug is mathematically sufficient to predict the treated response. A more detailed model that distinguishes between the two drugs’ mechanisms of action may provide more information about their synergy or competition.
@@ 2.3.3 Drug target identification
2.3.4 Uncertainty Quantification and Virtual Populations
A major advantage of QSP modeling is the ability to perform rigorous uncertainty quantification (UQ) to support regulatory decision-making. UQ is particularly valuable for evaluating benefit-risk profiles, justifying dose selection, and informing the design of confirmatory studies. However, when models are structurally unidentifiable, UQ analyses may become numerically unstable and computationally infeasible, limiting their utility. To overcome these limitations, we apply our model reduction workflow to generate a lower-dimensional, identifiable model amenable to Bayesian UQ and virtual population generation. This approach facilitates transparent characterization of variability in predicted outcomes and supports regulatory expectations for model credibility, including sensitivity to assumptions and relevance to the intended context of use.
We ask: given a representative response to untreated viral infection, can we characterize and quantify the range of responses to therapeutic intervention across a broader virtual population–prior to conducting a clinical trial? As a starting point, we use the parameter set calibrated to aggregated untreated data (Figure 4a) to represent the nominal disease trajectory. The model outputs of interest are viral load and T-cell time series under both untreated and bNAb-treated conditions. The untreated response serves as the reference data query, while the treated trajectory represents the target prediction. Our objective is to quantify uncertainty in these treatment outcomes and generate a virtual population reflecting plausible inter-individual variability in treatment response, thereby providing model-informed evidence to support benefit-risk assessments. This approach aligns with the regulatory emphasis on establishing a clear context of use for quantitative systems pharmacology models to inform drug development decisions and to support regulatory submissions by characterizing clinically relevant variability and informing trial design.
@@ 2.3.4 Uncertainty Quantification and Virtual Populations
MBAM gives a 16-parameter reduced model for these queries (Eq. (12) in the methods section). Since these queries only include viral load predictions, a subset of those used in section 2.3.2, the model is even simpler than before. Inspecting the resulting equations (see methods), six of the remaining parameters characterize the bNAb dynamics and therefore are trivially structurally unidentifiable from the untreated profile. For now, we hold these at their nominal values and consider the uncertainty in the remaining ten parameters. We then used this reduced model to generate a virtual population using Bayesian posterior sampling and categorized each virtual patient by its response both with and without treatment (see methods for more details).
Our response categorization builds on prior QSP analyses[95, 80] that showed that effective treatment leverages bistable dynamics in the model. We define successful treatment as shifting the equilibrium from the higher to the lower viral load steady state. Accordingly, we divide the virtual population into two broad groups: those whose parameter sets produce monostable dynamics (a single stable equilibrium) and those that exhibit bistability. The majority of virtual patients fall into the monostable category, meaning their dynamics converge to a single equilibrium that is not altered by treatment. While these patients may show a transient response, their long-term behavior remains unchanged, leaving little opportunity for long-term therapeutic benefit.
Bistable patients, in contrast, have two stable steady states: one corresponding to high viral load (progressive disease) and the other to low viral load (controlled disease). We further subdivide this group into four categories based on their response to treatment. Responder and nonresponder patients both converge to the high viral load state in the absence of treatment, but diverge under therapy: responders transition to the low viral load state with treatment, while nonresponders do not. Immune patients naturally reach the low viral load state without intervention, and treatment has no effect on their outcome. In contrast, adverse responders begin in the low viral load state without treatment, but therapy shifts their dynamics toward the high viral load state. Figure 7 illustrates representative dynamics for each of these five subpopulations.
Fig 7: Typical responses for each subpopulation. (A) The virtual population is first partitioned into patients whose dynamics exhibit monostability and those that exhibit bistability. Among the bistable patients, four subpopulations are identified as either (B) immune, i.e., low viral count both with and without treatment, (C) responders, i.e., high viral count without treatment, low viral count with treatment, (D) non responders, i.e., high viral count with and without treatment, and (E) adverse, i.e., low viral count without treatment, high viral count with treatment.
Table 1 summarizes the distribution of response categories within our 7,200-member virtual population. Notably, all five subpopulations are represented. However, only about 7% of patients are classified as responders who benefit from bNAb intervention. In contrast, fewer than 0.2% exhibit an adverse response, transitioning to higher viral load with treatment. The predicted steady state for most patients is unaffected by therapy, indicating that either they were monostable, i.e., having only a single steady state, or that the treatment did not perturb the dynamics sufficiently toward the other steady state, corresponding to either non-responders or immune.
Monostable Bistable
5683 1517
Responder Nonresponder Immune Adverse
486 576 445 10
Table 1: Categorization of the 7200-member virtual population generated from Bayesian posterior sampling. Patients are first separated by monostable or bistable dynamics. Bistable patients are further separated by response both with and without bNAb therapy.
The statistics of the virtual population, such as the relative size of each subpopulation, are determined by the prior distribution used in the Bayesian sampling process. For this study, we employed a flat prior over the logarithm of the model parameters, centered on the nominal values obtained by fitting to aggregated untreated data. The prior spans 16 orders of magnitude in log-space, corresponding to a fold-change of nearly 10 million. For this study, we have selected a prior that is intentionally broad to sample the full range of possible responses. Roughly speaking, the relative distribution of the subpopulations reflects the volume of parameter space corresponding to each category. As such, it does not necessarily reflect the relative distribution of responses in real patients whose parameters are unlikely to be uniformly distributed, nor so broad. Importantly, the Bayesian framework applied to the reduced model allows us to sample efficiently from the posterior for any given prior, ensuring that all assumptions are explicitly defined at the outset. In some applications, prior information about the natural distribution of parameters or responses could be used to construct different virtual populations. This transparency represents a significant step forward in the standardization and reproducibility of virtual population generation.
Because the virtual population is derived from a minimal model, we can directly identify which parameters determine the category to which a particular virtual patient belongs. This, in turn, enables mechanistic reasoning about the origins of divergent responses and can inform strategies for monitoring safety during clinical trials or treatment stratification. Using a random forest classifier, we find that the parameters $\xi$, $d_{E}$, and $d_{Q}$ are most strongly associated with the separation between monostable and bistable patients, as shown in Figure 8a. This suggests that the bifurcation separatrix is most sensitive to variation along these parameter directions. A 10% cross-validation of this model could accurately predict mono- vs bistability for 94.5% of out-of-training-set simulations ($\kappa=0.83$).
Fig 8: Subpopulations and various parameter values. (A) A random forest classifier identifies the parameters $\xi$, $d_{E}$, and $d_{Q}$ as the strongest predictor for whether a patient will be monostable or bistable. (B) The parameters $d_{T}$, $p_{I}$, $c$, and $d_{Q}$ are most useful for predicting whether a bistable patient will be a responder, non-responder, immune, or adverse.
Among the bistable patients, the parameters $d_{T}$, $p_{I}$, $c$ and $d_{Q}$ were most predictive of subpopulation membership (Figure 8b). Cross-validation of a random forest classifier trained on these parameters yielded 89% accuracy in predicting bistable subtyptes ($\kappa=0.83$). Due to the extremely low prevalence of adverse responders in our virtual population (0.13%), there were insufficient data to robustly identify the specific parameter combinations driving adverse outcomes. Nevertheless, qualitative analysis reveals that the basin of attraction associated with immune control is consistently narrow in these individuals. This narrow basin explains both the rarity of adverse responders and the susceptibility of their trajectories to be perturbed into the progressive disease state. These findings suggest that transient therapeutic perturbations—such as CART therapy or transient effector cell suppression—could shift system dynamics across the separatrix and redirect trajectories toward the healthier equilibrium.
From a clinical and regulatory perspective, the virtual population also supports safety evaluation and early biomarker identification. Across all responder patients, CD4+ T-cell counts remained stable at over 90% of uninfected levels over the full 1,000-day simulation period. In contrast, other subpopulations—particularly non-responders and adverse responders—exhibited rapid T-cell decline to 10–80% of baseline levels, often within the first 30 days post-infection. These model-based predictions highlight the utility of early T-cell monitoring as a prognostic biomarker. Specifically, CD4+ T-cell counts during the initial month post-infection may serve as a reliable surrogate endpoint for evaluating long-term therapeutic efficacy of bNAb-based interventions.